acities at constant pressure for O(g), Odg) and Olg) are 21.90, 29.36 and 39.20 I K 3mol, respectively. Discuss the difference using the equipartition theorem.
Answer #1

According to the equipartition of energy (theorem), every degree of freedom has an average energy of rac{1}{2}k_{B}T in thermal equilibrium and contributes  rac{1}{2}k_{B} to the system's heat capacity. Here k_{B} is Boltzmann constant and T is absolute temperature.

i.e. if degree of freedom of an atom is 'f' then we can write total energy of the atom as

E = rac{f}{2}k_{B}T

And if there is one moles of the atom then

  E = rac{f}{2}RT ..................................(1)

Where R = Nk_{B} is universal gas constant and N is Avogadro constant.

The molar heat capacity at constant volume (C_{v}), is defined as

C, s (%) ..............................(2)

or,   C, RT

or, C_{v} = rac{f}{2} R ................................(3)

And the relation between molar heat capacity at constant volume  (C_{v}) and the molar heat capacity at constant pressure is

  C_{p} - C_{v} = R

or, C_{p} = C_{v} + R ..............................(4)

Now for oxygen atom (i.e. O) :

As oxygen atom is monoatomic, so it will have only 3 translational degrees of freedom. i.e. f = 3.

So.

c,--R-1.5 × 8.314 12.471 JK-1mol-1

Therefore C,-C, + R = 12.471+ 8.314 20.785 JK-1 mol-1

For oxygen gas (i.e. O2) :
Total degree of freedom of O 2 (diatomic molecule) is 5 (3 translation + 2 rotation) i.e f = 5 .
So, C.--R = 2.5 × 8.314 = 20.785 JK-1 mol-1 G, =
Therefore c, = c, + R-20.785 + 8.314 29.099 JK-1 mol-1
For ozone gas (i.e. O3) :

Total degree of freedom of O3 (triatomic molecule) is 7 (3 translation + 3 rotation + 1 vibration) i.e f = 7 .

So,   7 c,ー-R = 3.5 × 8.314 29.099 JK-1mol-1

Therefore   c,-C,, + R = 29.099 + 8.314-37.413 JK-mol-1

So, according to equipartition theorem the heat capacity at constant pressure for O (g) , O2 (g) and O3 (g) is 20.785 , 29.099 and 37.413 J/K-mol respectively.

Hence, we can see that these calculated values differ from originally given values (which are 21.90 , 29.36 and 39.20 J/K-mol) by 1 to 2 units.

For any doubt please comment and please give an up vote.